# Gauge preservation in renormalization for Yang-Mills and gravity theories

### v2

@@ Abstract
In the present work, multiplicative renormalization [1] for Yang-Mills theories is reviewed. While this subject is not new, it is suggested that a clear understanding of these methods leads to a systematic way for interpreting the counter terms needed for non multiplicative renormalization of quantum gravity, for models such as [2]-[6]. These models are renormalizable but contain apparent instabilities leading to possible unitarity loss, an earlier analog is [7]. This systematic may be interesting, especially in the modern context, since there are efforts for avoiding with those apparent instabilities by employing variants of the standard quantization methods [8]-[22].
@@ 1. Introduction
The discovery of the renormalizability property of Yang-Mills theories [23]-[25] marked a new era in QFT and particle physics and, in particular, leaded to the development of the Standard Model. There are of course several aspects about confining theories which are still not understood. However, the consensus is that Yang-Mills theories are correct and that the methods still available are not suited for studying their predictions in full detail.
Gravity theories on the other hand are more elusive, and the dream of constructing a renormalizable and unitary model of gravitation eluded success for years. There are models that are close to achieve several features that a consistent gravity model has to have, however, there is always a missing point which invalidate the model or request further improvements. An example are the quadratic models of gravity, which posses fourth order equations of motion and were considered as a plausible quantum gravity candidates in [4]-[4]. These models were studied further by Stelle in [5]-[6], in the context of modern QFT. These classic references, in particular [6] consider the path integral quantization of theories, the problem of gauge fixing and the solution of the Slavnov-Taylor identities for these scenarios.
The authors of [4]-[6] acknowledge however, that these theories apparently posses a ghost in the physical spectrum. This suggest instabilities and unitarity loss [7]. The point is propagator of such four order theory will be given in a Pauli Villars form as
@@
For large values of $k$ this propagator will go as $M^{2}k^{-4}$. In QFT without gravity, the mass $M$ is interpreted as a cutoff, which tends $M\to\infty$ at the end of calculations. In the gravity approach, this cutoff is rendered finite, and the minus sign in the last propagator expression suggests the presence of an state with negative kinetic energy, which may indicate a potential instability.
Even though this drawback, it should be mentioned that this type of view about instabilities have been argued in some modern references. An incomplete list is [8]-[21]. These references suggest that the problem of the apparent instabilities may be solved by employing variants of the quantization method, some of them inspired in earlier works of Pauli and Dirac. This revivals the interest in those quadratic models since, if the ghost problem is overcome, they will become a very promising candidate for a quantum gravity model.
@@
The reader may notice that the present exposition lacks details in some aspects and is pretty detailed in some others. The details that are omitted can be found in [36]-[41] or [27]-[31], and it is not the intention to compete with those clear expositions. The detailed parts of the present text correspond to situations which in our opinion are not discussed in the available literature and are potentially useful. The idea is that the reader may complement the literature with the present details in order to get some new insights. It may be said that the text contains the explanations that we would have liked to have when learning the subject. The reward of the effort is that an interpretation about the counter terms in the gravity scenario appears that, at the best of the authors knowledge, is not particularly described in the available literature. Recent modern literature dealing with topics of interest such as the background field method or renormalization group may be found in [32]-[35].
The present work is briefly organized as follows. Section 2 discusses the Slavnov-Taylor identities in the context of Yang-Mills theories, and contains detailed explanations about how multilplicative renormalization is achieved and how it insures BRST invariance at a given order in $\hbar$. Section 3 contains the arguments about the renormalizability of quadratic gravity, and the explicit solutions of the Slavnov-Taylor identities for these models. The solution is interpreted mantaining some analogy with multiplicative renormalization. Even in the gravity case the renormalization is not multiplicative, it is shown that BRST invariance is preserved order by order by the addition of non standard coupling to the initial action, which still respect BRST symmetry. The last section contains some possible applications.
@@ 2.1 Preliminars about Yang-Mills theories
$G=e^{-iT_{a}\theta^{a}},\qquad a=1,.,\text{dim}\;G.$
@@
At infinitesimal leve the gauge transformation (2.1) is given by
@@ 2.2 Peculiarities about gauge fixing and quantization
This integral sum a fixed configuration $A^{af}_{\mu}(x)$ and all its gauge transformed representatives $A^{afG}_{\mu}(x)$. As there is a plethora of such transformed representatives, these representatives contribution leads to a divergent value of $Z(J)$. The cure for this problem is to impose a gauge fixing, which is a procedure allowing to express the divergent part of $Z(J)$ as an overall divergent factor that can be discarded. Leaving aside issues such as the Gribov ambiguities [45], which may obstruct the standard gauge fixing procedure, in the following, linear gauges of the form $F^{a}(A^{b})=0$ with
@@
In a non abelian theory, a gauge transformation (2.1) is induced by a group element, $G$ corresponding to parameters $\theta^{a}$ introduced in that formula. Under a small gauge transformation it is expected that
@@
$\frac{\delta F}{\delta\theta(y)}=\partial_{\mu}\partial^{\mu}\delta^{4}(x-y).$
@@
$M(x,y)=\partial_{\mu}\partial^{\mu}\delta^{4}(x-y).$
@@
where the Faddev-Popov factor $\Delta_{FP}(A^{b})$ has been introduced. The Faddev-Popov determinant is whatever is needed for making the last expression equal to one , this ensures that the insertion of the last factor inside $Z(J)$ will not change its value. The Dirac delta dictates the choice of gauge, as depends on $F^{a}(A_{\mu}^{b})$ explicitly. On the other hand, the formulas (2.12)- and (2.12) imply that this factor may be represented in terms of ghost and anti-ghost as
@@
Due to the gauge invariance property $\Delta_{FP}(A^{bG}(x))=\Delta_{FP}(A^{b}(x))$ it follows that the last expression is
$Z(0)=\int DG\int DA_{\mu}e^{iS_{y}}\Delta_{FP}(A^{b}(x)).$ (2.15)
as the Yang-Mills action $S_{y}$ is also gauge invariant [53]. The integration on $DG$ is an overall infinite factor that can be discarded. By this, and taking into account the representation (2.13) for the Faddev-Popov determinant, it is arrived to the gauge fixed functional
[…]
@@
As a final comment, it should be recalled that at the classical level for Yang-Mills theories, it is customary to introduce a gauge fixing term inside the lagrangian as follows
$L_{gf}=L_{y}-\frac{\lambda}{2}F_{a}(A^{b}_{\mu})^{2}.$
This gauge fixing term can be introduced into the action (2.16) by a change of the normalization inside the path integrals, as specified in pages 249-250 of [41]. This means that the alternative expression
$Z(J)=\int DA^{b}_{\mu}D\overline{\eta}^{a}D\eta^{c}\exp\{iS_{y}-\frac{i\lambda}{2}\int d^{4}xF_{a}(A^{b}_{\mu})^{2}\}$ (2.17)
$\exp\{-i\int d^{4}yd^{4}x\,(\overline{\eta}^{a}(x)M_{ac}(x,y)\eta^{c}(y))+i\int d^{4}x\;J_{\mu}A^{\mu}\},$
may be used as well for the path integral. This expression will be important in the next sections.
@@ 2.3 Generalizations of gauge symmetry
The expression (2.17) for the partition function leads to the gauge fixed action
$S_{gf}=\int\;d^{4}x\;[L_{y}-\frac{\lambda}{2}F_{a}(A^{b}_{\mu})^{2}-\overline{\eta}^{a}M_{ab}\eta^{b}],$ (2.18)
which, of course, is non gauge invariant, as it reflects an specific choice of gauge. An important task is to determine if there is a generalization of the gauge symmetry (2.2), involving $A^{b}_{\mu}$, $\eta^{a}$ and $\overline{\eta}^{a}$ leaving $S_{gf}$ invariant. Denote this potential new symmetry as $\delta_{n}$, where the subscript $n$ emphasizes its novelty. Since the component $L_{y}$ involves only $A_{\mu}$ and is gauge invariant, it may be natural, although not necessarily true, to expect that $\delta_{n}A^{a}_{\mu}=D_{\mu}^{ab}\theta^{b}$. By employing this ansatz, the goal is to determine how $\delta_{n}$ acts on the remaining fields $\eta^{a}$ and $\overline{\eta}^{a}$. This has to be determined with the requirement that the variation of the remaining terms in the gauge-fixed action,
$L_{g}=-\frac{\lambda}{2}F_{a}(A_{\mu}^{b})^{2}-\overline{\eta}^{a}M_{ab}\eta^{b}$ (2.19)
@@
The transformations (2.21)-(2.22) or (2.23) given above still does not describe a symmetry, unless $\delta(M_{ab}\eta^{b})=0$. Since these transformations (2.21)-(2.22) and (2.23) affect the anti-ghost field $\overline{\eta}^{a}$ and the gauge field $A_{\mu}^{a}$, the requirement $\delta(M_{ab}\eta^{b})=0$ has to be employed to fix the transformation of the ghost $\eta^{a}$, which is still missing. An attempt could be $\delta_{n}\eta^{a}=0$. The issue is that, in a generic situation, $M_{ab}$ contains the gauge field $A^{a}_{\mu}$ inside the covariant derivative $D_{\mu}^{cb}$, rendering this ansatz inefficient. This problem, of course, is avoided if $c_{abc}=0$, which corresponds to abelian models. In this case $D_{\mu}^{ab}$ reduces to $\partial_{\mu}$, which is obviously $A^{\mu}$ independent. This implies that, for abelian models, a generalization of the gauge symmetry is described by
@@
where the anti commutation property of $\eta^{a}\eta^{b}=-\eta^{b}\eta^{a}$ was employed in order to delete the derivative $\partial_{\mu}$ and consequently obtain this relation. However, the terms that do not contain $\partial_{\mu}$ has to cancel each other as well. Fortunately they do, due to the Jacobi identity. To see this, observe that the remaining terms not proportional to $\partial_{\mu}$, denoted with the letter $O$ of obstruction in the following
@@
In order to take the ghost and gauge fields as common factor, this may be written as
@@
A final important comment is that given a function $F(A^{a}_{\mu},\eta^{a})$, then $s^{2}F=0$ due to the above derived nilpotency conditions. In fact
$\delta F=\frac{\partial F}{\partial A_{\mu}^{a}}sA_{\mu}^{a}\epsilon+\frac{\partial F}{\partial\eta^{a}}s\eta^{a}\epsilon,$
@@
$\delta sF=\frac{\partial F}{\partial A_{\mu}^{a}\partial A_{\nu}^{b}}\delta A_{\nu}^{b}sA_{\mu}^{a}+\frac{\partial F}{\partial\eta^{a}\partial\eta^{b}}s\eta^{b}\epsilon s\eta^{a}+\frac{\partial F}{\partial A_{\mu}^{a}\partial\eta^{b}}(s\eta^{b}\epsilon sA_{\mu}^{a}+sA_{\mu}^{a}\epsilon s\eta^{b})$
@@ 2.4 Connected Green functions and generalizations of BRST symmetry
$\eta^{a}\to e^{i\alpha}\eta^{a},\qquad\overline{\eta}^{a}\to e^{-i\alpha}\overline{\eta}^{a},\qquad\xi_{a}\to e^{-i\alpha},\qquad\overline{\xi}_{a}\to e^{i\alpha}\overline{\xi}_{a},$ (2.36)
@@
From (2.35), it is seen that the currents $K_{\mu}^{a}$ and $L_{a}$ have mass dimension two, so they may appear at most in a quadratic term. Thus, they may appear at most in terms such as
@@
$L_{L}=L_{a}\partial^{\mu}A_{\mu}^{a}$
is Lorenz invariant and has the correct mass dimensions, however, it does not have ghost number zero. A term of the form
@@
The second important question to answer is, which is the most general symmetry $s_{n}$ generalizing (2.30) increasing the ghost number by one, and which is nilpotent when acting on $A_{\mu}^{a}$ and $\eta^{a}$. The condition of increasing the ghost number by one is because it is evident that the standard BRST symmetry (2.30) has precisely this property. Dimensional analysis and ghost increase number requirement indicate that
$s_{n}A_{a}^{\mu}=a_{ab}\partial_{\mu}\eta^{b}-gd_{abc}A^{b}_{\mu}\eta^{c},\qquad s_{n}\eta^{a}=-\frac{g}{2}e_{abc}\eta^{b}\eta^{c}.$ (2.39)
@@
$\delta_{\epsilon}(s_{n}A_{a}^{\mu})=a_{ab}\partial_{\mu}\delta_{\epsilon}\eta^{b}-gd_{abc}\delta_{\epsilon}(A^{b}_{\mu})\eta^{c}-gd_{abc}A^{b}_{\mu}\delta_{\epsilon}(\eta^{c})$
@@
while the vanishing of the terms proportional to $g^{2}$ leads to
@@
By interpreting $e_{bec}$ as structure constants of a Lie algebra, it follows that $T^{a}$ lies in the adjoint representation. In other words $d_{abc}=e_{abc}$. Finally, the fact that $d_{abc}a_{bd}=0$ implies that $a_{ab}$ commutes with all the generators of the algebra. If the algebra is simple, that means that $a_{ab}=a\delta_{ab}$. This will be the case to be considered below. Therefore, the generalization of BRST symmetry for simple Lie algebras reads as follows
$s_{n}A_{a}^{\mu}=a\partial_{\mu}\eta^{a}-gd_{abc}A^{b}_{\mu}\eta^{c},\qquad s_{n}\eta^{a}=-\frac{g}{2}d_{abc}\eta^{b}\eta^{c},\qquad s_{n}\overline{\eta}^{a}=\lambda F^{a}(A_{\mu}^{b}).$ (2.40)
@@
The third and last problem is to classify the most general action $S_{nbrst}(A,\eta,\overline{\eta},K,L)$ which has zero ghost number, it is invariant under Lorenz transformations, has dimension four or less due to renormalizability, and depends on the fields $K^{\mu}_{a}$ and $\eta_{a}$ only through the linear combination $K^{\mu}_{a}-(\overline{\eta}\phi^{\mu})_{a}$. The main difference with the above analysis is that BRST invariance is not required. Although this last question may seem less motivated than the above ones, this analysis will be useful when dealing with Slavnov-Taylor identities. As the combination $K^{\mu}_{a}-(\overline{\eta}\phi^{\mu})_{a}$ has ghost number $-1$ due to (2.37) and mass dimension $2$ due to (2.35) the only way that it may appear is
@@
In order to close this section, it should be mentioned that the references [36]-[37] study similar questions, with the difference that they use an enlarged version of the BRST symmetry with new auxiliary fields, nilpotent for all of them. The action is expressed as the sum of the Yang-Mills one and a pure BRST term $sF$ with $F$ some function of the fields. The BRST invariance is easier to be proved with this enlarged field content. Furthermore, those expositions are adapted to more general gauges, and this is of clear interest. The present approach is minimalist, and for this reason this general and important formulation is skipped.

### v3

@@ Abstract
In the present work, multiplicative renormalization [1] for Yang-Mills theories is reviewed. While this subject is not new, it is suggested that a clear understanding of these methods leads to a systematic way for interpreting the counter terms needed for non multiplicative renormalization of quantum gravity, for models such as the ones introduced in [2]-[6]. These models are renormalizable, however they contain apparent instabilities leading to possible unitarity loss, an earlier analog is [7]. The systematic procedure discussed along the text may be interesting, especially in the modern context, since there are efforts for avoiding with those apparent instabilities by employing variants of the standard quantization methods [8]-[22].
@@ 1. Introduction
The discovery of the renormalizability of Yang-Mills theories [23]-[25] marked a new era in QFT and particle physics and, in particular, led to the development of the Standard Model. There are several aspects about confining theories which are still not understood. The consensus is that Yang-Mills theories are correct and that the methods still available are not suited for studying their predictions in full detail.
Gravity theories on the other hand are more elusive, and the dream of constructing a renormalizable and unitary model of gravitation eluded success for years. There are models close to achieve several features that a consistent gravity model has to have, however, there is always a missing point which invalidate the model or request further improvements. Quadratic gravity models are examples, these scenarios posses fourth order equations of motion and were considered as a plausible quantum gravity candidates in [2]-[4]. Their properties were studied further by Stelle in [5]-[6], in the context of modern QFT. These classic references, in particular [6] consider the path integral quantization of theories, the problem of gauge fixing and the solution of the Slavnov-Taylor identities for these scenarios.
The authors of [4]-[6] recognize however, that these theories apparently posses a ghost in the physical spectrum. This suggest instabilities and unitary loss [7]. The point is propagator of such four order theory will be given in a Pauli Villars form as
@@
For large values of $k$ this propagator will go as $M^{2}k^{-4}$. In QFT without gravity, the mass $M$ is interpreted as a cutoff, which tends $M\to\infty$ at the end of calculations. In the gravity approach, this cutoff is rendered finite, and the minus sign in the last propagator expression suggests the presence of a state with negative kinetic energy, which may indicate a potential instability.
Even taking into account this apparent drawback, it should be mentioned that this type of view about instabilities have been discussed in some modern references. An incomplete list is [8]-[21]. These references suggest that the problem of the apparent instabilities may be solved by employing variants of the quantization method, some of them inspired in earlier works of Pauli and Dirac. This revivals the interest in those quadratic models since, if the ghost problem is overcome, they will become a very promising candidate for a quantum gravity model.
@@
The reader will probably notice that the present exposition lacks details in some aspects and is pretty detailed in some others. The details that are omitted can be found in [36]-[41] or [27]-[31], and it is not the intention to compete with those clear expositions. The detailed parts of the present text correspond to situations which in our opinion are not discussed in the available literature and are potentially useful. The idea is that the reader may complement the literature with the present details in order to get some new insights. It may be said that the text contains the explanations that we would have liked to have when learning the subject. The reward of the effort is that an interpretation about the counter terms in the gravity scenario appears that, at the best of the authors knowledge, is not particularly described in the available literature. Recent modern literature dealing with topics of interest such as the background field method or renormalization group may be found in [32]-[35].
The present work is briefly organized as follows. Section 2 discusses the Slavnov-Taylor identities in the context of Yang-Mills theories, and contains detailed explanations about how multiplicative renormalization is achieved and how it insures BRST invariance at a given order in $\hbar$. Section 3 contains the arguments about the renormalizability of quadratic gravity, and the explicit solutions of the Slavnov-Taylor identities for these models. The solution is interpreted by keeping some analogy with multiplicative renormalization. Even in the gravity case the renormalization is not multiplicative, it is shown that BRST invariance is preserved order by order by the addition of non standard coupling to the initial action, which still respect BRST symmetry. The last section contains some possible applications.
@@ 2.1 Preliminars about Yang-Mills theories
$G=e^{-iT_{a}\theta^{a}},\qquad a=1,.,\text{dim}\;G,$
@@
At infinitesimal level the gauge transformation (2.1) may be expressed as
@@ 2.2 Peculiarities about gauge fixing and quantization
This integral sum a fixed configuration $A^{af}_{\mu}(x)$ and all its gauge transformed representatives $A^{afG}_{\mu}(x)$. However, the integral one would like to calculate is
$Z_{g}(J)=\int[DA_{\mu}]e^{i[S_{y}+i\int d^{4}xJ^{\mu}A^{a}_{\mu}]},$ (2.5)
with the bracket [][][ ] indicating one representative of a full gauge transformed configuration. The resulting measure $[DA_{\mu}]$, as it is well known and will be described below, contains non trivial factors. As there is a plethora of such transformed representatives, their contribution leads to a divergent value of $Z(J)$. The cure for this problem is to impose a gauge fixing, which is a procedure allowing to express the divergent part of $Z(J)$ as an overall divergent factor that can be discarded. Explicitly, it is expected that
$Z(J)=Vol(G)Z_{g}(J).$
Leaving aside issues such as the Gribov ambiguities [45], which may obstruct the standard gauge fixing procedure, in the following, linear gauges of the form $F^{a}(A^{b})=0$ with
@@
In a non abelian theory, a gauge transformation (2.1) is induced by a group element, $G$ corresponding to parameters $\theta^{a}$ introduced in that formula. Under a small gauge transformation it generally given as
@@
$\frac{\delta F}{\delta\theta(y)}=\partial_{\mu}\partial^{\mu}\delta^{4}(x-y),$
@@
$M(x,y)=\partial_{\mu}\partial^{\mu}\delta^{4}(x-y).$ (2.9)
@@
where the Faddev-Popov factor $\Delta_{FP}(A^{b})$ has been introduced. The Faddev-Popov determinant is whatever is needed for making the last expression equal to one , this ensures that the insertion of the last factor inside $Z(J)$ will not change its value. The Dirac delta dictates the choice of gauge, as depends on $F^{a}(A_{\mu}^{b})$ explicitly. On the other hand, the formulas (2.13) and (2.14) imply that this factor may be represented in terms of ghost and anti-ghost as
@@
The next task is to eliminate the Delta function, which has the gauge invariance property $\Delta_{FP}(A^{bG}(x))=\Delta_{FP}(A^{b}(x))$. This procedure is described in pages 228-237 in the book [28], see also [27], [41]. By taking into account the representation (2.15) a for the Faddev-Popov determinant, the resulting gauge fixed functional is
$Z(J)=\int DA^{b}_{\mu}D\overline{\eta}^{a}D\eta^{c}\exp\{iS_{y}-\frac{i\lambda}{2}\int d^{4}xF_{a}(A^{b}_{\mu})^{2}\}$
$\exp\{-i\int d^{4}yd^{4}x\,(\overline{\eta}^{a}(x)M_{ac}(x,y)\eta^{c}(y))+i\int d^{4}x\;J_{\mu}A^{\mu}\}.$ (2.17)
The action in the last expression corresponds to a lagrangian which contains the gauge fixing term
$L_{gf}=L_{y}-\frac{\lambda}{2}F_{a}(A^{b}_{\mu})^{2}.$
The terms proportional to $M_{ab}(x,y)$ correspond to the Faddeev-Popov determinant. This expression will be important in the next sections.
@@ 2.3 Generalizations of gauge symmetry
The expression (2.17) for the Feynman path integral leads to the gauge fixed action
$S_{gf}=\int\;d^{4}x\;[L_{y}-\frac{\lambda}{2}F_{a}(A^{b}_{\mu})^{2}-\overline{\eta}^{a}M_{ab}\eta^{b}].$ (2.18)
Here the operator $M_{ab}(x,y)$ usually has a delta such as in (2.9) has been converted in one depending on one variable $M_{ab}(x)$, in analogous fashion as in (2.11). The last action, of course, is non gauge invariant, as it reflects an specific choice of gauge. An important task is to determine if there is a generalization of the gauge symmetry (2.2), involving $A^{b}_{\mu}$, $\eta^{a}$ and $\overline{\eta}^{a}$ leaving $S_{gf}$ invariant. Denote this potential new symmetry as $\delta_{n}$, where the subscript $n$ emphasizes its novelty. Since the component $L_{y}$ involves only $A_{\mu}$ and is gauge invariant, it may be natural, although not necessarily true, to expect that $\delta_{n}A^{a}_{\mu}=D_{\mu}^{ab}\theta^{b}$. By employing this ansatz, the goal is to determine how $\delta_{n}$ acts on the remaining fields $\eta^{a}$ and $\overline{\eta}^{a}$. This has to be determined with the requirement that the variation of the remaining terms in the gauge-fixed action,
$L_{g}=-\frac{\lambda}{2}F_{a}(A_{\mu}^{b})^{2}-\overline{\eta}^{a}M_{ab}\eta^{b},$ (2.19)
@@
The transformations (2.21)-(2.22) or (2.23) given above still do not describe a symmetry, unless $\delta(M_{ab}\eta^{b})=0$. Since these transformations (2.21)-(2.22) and (2.23) affect the anti-ghost field $\overline{\eta}^{a}$ and the gauge field $A_{\mu}^{a}$, the requirement $\delta(M_{ab}\eta^{b})=0$ has to be employed to fix the transformation of the ghost $\eta^{a}$, which is still missing. An attempt could be $\delta_{n}\eta^{a}=0$. The issue is that, in a generic situation, $M_{ab}$ contains the gauge field $A^{a}_{\mu}$ inside the covariant derivative $D_{\mu}^{cb}$, rendering this ansatz inefficient. This problem, of course, is avoided if $c_{abc}=0$, which corresponds to abelian models. In this case $D_{\mu}^{ab}$ reduces to $\partial_{\mu}$, which is obviously $A^{\mu}$ independent. This implies that, for abelian models, a generalization of the gauge symmetry is described by
@@
where the anti commutation property of $\eta^{a}\eta^{b}=-\eta^{b}\eta^{a}$ was employed in order to delete the derivative $\partial_{\mu}$ and consequently obtaining this relation. However, the terms that do not contain $\partial_{\mu}$ has to cancel each other as well. Fortunately they do, due to the Jacobi identity. To see this, observe that the remaining terms not proportional to $\partial_{\mu}$, denoted with the letter $O$ of obstruction in the following
@@
In order to take the ghost and gauge fields as common factor, this may be expressed in the following form
@@
A final important comment is that given a function $F(A^{a}_{\mu},\eta^{b})$, then $s^{2}F=0$ due to the above derived nilpotency conditions. In fact
$\delta F=\frac{\partial F}{\partial A_{\mu}^{a}}sA_{\mu}^{a}\epsilon+\frac{\partial F}{\partial\eta^{b}}s\eta^{b}\epsilon,$
@@
$\delta sF=\frac{\partial F}{\partial A_{\mu}^{a}\partial A_{\nu}^{b}}\delta A_{\nu}^{b}sA_{\mu}^{a}+\frac{\partial F}{\partial\eta^{a}\partial\eta^{b}}s\eta^{b}\epsilon s\eta^{a}+\frac{\partial F}{\partial A_{\mu}^{a}\partial\eta^{b}}(s\eta^{b}\epsilon sA_{\mu}^{a}+sA_{\mu}^{a}\epsilon s\eta^{b}).$
@@ 2.4 Connected Green functions and generalizations of BRST symmetry
$\eta^{a}\to e^{i\alpha}\eta^{a},\qquad\overline{\eta}^{a}\to e^{-i\alpha}\overline{\eta}^{a},\qquad\xi_{a}\to e^{-i\alpha}\xi_{a},\qquad\overline{\xi}_{a}\to e^{i\alpha}\overline{\xi}_{a},$ (2.36)
@@
First problem: From (2.35), it is seen that the currents $K_{\mu}^{a}$ and $L_{a}$ have mass dimension two, so they may appear at most in a quadratic term. Thus, they may appear at most in terms such as
@@
$L_{L}=L_{a}\partial^{\mu}A_{\mu}^{a},$
is Lorenz invariant and has the correct mass dimensions, however, it does not have ghost number zero. A term of the following type
@@
Second problem: The second important question to answer is, which is the most general symmetry $s_{n}$ generalizing (2.30) increasing the ghost number by one, and which is nilpotent when acting on $A_{\mu}^{a}$ and $\eta^{a}$. The condition of increasing the ghost number by one is because it is evident that the standard BRST symmetry (2.30) has precisely this property. Dimensional analysis and ghost increase number requirement indicate that
$s_{n}A^{a}_{\mu}=a_{ab}\partial_{\mu}\eta^{b}-gd_{abc}A^{b}_{\mu}\eta^{c},\qquad s_{n}\eta^{a}=-\frac{g}{2}e_{abc}\eta^{b}\eta^{c}.$ (2.39)
@@
$\delta_{\epsilon}(s_{n}A^{a}_{\mu})=a_{ab}\partial_{\mu}\delta_{\epsilon}\eta^{b}-gd_{abc}\delta_{\epsilon}(A^{b}_{\mu})\eta^{c}-gd_{abc}A^{b}_{\mu}\delta_{\epsilon}(\eta^{c})$
@@
while the vanishing of the terms proportional to $g^{2}$ results in the following relation
@@
By interpreting $e_{bec}$ as structure constants of a Lie algebra, it follows that $T^{a}$ lies in its adjoint representation. In other words $d_{abc}=e_{abc}$. Finally, the fact that $d_{abc}a_{bd}=0$ implies that $a_{ab}$ commutes with all the generators of the algebra. If the algebra is simple, that means that $a_{ab}=a\delta_{ab}$. This will be the case to be considered below. Therefore, the generalization of BRST symmetry for simple Lie algebras reads as follows
$s_{n}A^{a}_{\mu}=a\partial_{\mu}\eta^{a}-gd_{abc}A^{b}_{\mu}\eta^{c},\qquad s_{n}\eta^{a}=-\frac{g}{2}d_{abc}\eta^{b}\eta^{c},\qquad s_{n}\overline{\eta}^{a}=\lambda F^{a}(A_{\mu}^{b}).$ (2.40)
@@
The third problem: The third and last problem is to classify the most general action $S_{nbrst}(A,\eta,\overline{\eta},K,L)$ which has zero ghost number, it is invariant under Lorenz transformations, has dimension four or less due to renormalizability, and depends on the fields $K^{\mu}_{a}$ and $\eta_{a}$ only through the linear combination $K^{\mu}_{a}-(\overline{\eta}\phi^{\mu})_{a}$. The main difference with the above analysis is that BRST invariance is not required. Although this last question may seem less motivated than the above ones, this analysis will be useful when dealing with Slavnov-Taylor identities. As the combination $K^{\mu}_{a}-(\overline{\eta}\phi^{\mu})_{a}$ has ghost number $-1$ due to (2.37) and mass dimension $2$ due to (2.35) the only way that it may appear is
@@
In order to close this section, it should be mentioned that the references [36]-[37] study similar questions, with the difference that they use an enlarged version of the BRST symmetry with new auxiliary fields, nilpotent for all of them. These references employ the Nakanishi-Lauptrup formalism [61]-[62] . The action is expressed as the sum of the Yang-Mills one and a pure BRST term $sF$ with $F$ some function of the fields. The BRST invariance is easier to be proved with this enlarged field content. Furthermore, those expositions are adapted to more general gauges, and this is of clear interest. Another formalism of particular interest is the Batalin-Vilkovisky one [63]-[64], which is valid even for open symmetry algebras. The present approach is minimalist, and for this reason this general and important formulation is skipped.